# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v3

@@ Appendix C Derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$
In this section, we show the derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$ for $\rho=0$. This result is obtained directly from [59] as also elaborated in [57, Chap. 1]. Let $\tilde{f}\left(\rho,S_{1}\right)\triangleq\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}=\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}$, $A=\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}=\exp{\log\left({\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}}\right)}$. Thus,
$\frac{\partial}{\partial\rho}\left(A\right)=A\frac{\partial}{\partial\rho}\left(\log\left(A\right)\right)=A\frac{\partial}{\partial\rho}\left(\left(1+\rho\right)\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)\right)$
@@
=$\overset{(b)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))$
$\hskip 17.07182pt-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
=$\overset{(c)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))$
$\hskip 17.07182pt-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
=$\overset{(d)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}$
$\hskip 17.07182ptP_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
@@
$\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}$
$\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
@@
$=$ $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)$
$\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)}{P_{W}\left(S_{1},S_{2}\right)}\Biggr{)}$

### v4

@@ Appendix C Derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$
In this section, we show the derivative of $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$ for $\rho=0$. This result is obtained directly from [62] as also elaborated in [42, Chap. 1]. Let $\tilde{f}\left(\rho,S_{1}\right)\triangleq\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}=\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}$, $A=\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}=\exp{\log\left({\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}}\right)}$. Thus,
$\frac{\partial}{\partial\rho}\left(A\right)=A\frac{\partial}{\partial\rho}\left(\log\left(A\right)\right)$
$=A\frac{\partial}{\partial\rho}\left(\left(1+\rho\right)\log\left(\tilde{f}\left(\rho,S_{1}\right)\right)\right)$
@@
=$\overset{(b)}{=}\tilde{f}\left(\rho,S_{1}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
=$\overset{(c)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{1+\rho}\Biggr{(}\log(\tilde{f}\left(\rho,S_{1}\right))-\frac{1}{(1+\rho)}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\left(P_{W}\left(S_{1},S_{2}\right)\right)\Biggr{)}$
=$\overset{(d)}{=}\left(\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
@@
$\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(S_{1},S_{2}\right)^{\frac{1}{\left(1+\rho\right)}}}\Biggr{)}.$
@@
$=$ $\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}\Biggr{(}\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)\log\frac{\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(S_{1},S_{2}\right)}{P_{W}\left(S_{1},S_{2}\right)}\Biggr{)}$